SpaceRef

SpaceRef


NASA Space Station On-Orbit Status 6 April 2005

Status Report From: NASA HQ/SpaceRef/NASA Watch
Posted: Wednesday, April 6, 2005

image

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2005) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.

After reconfiguring the EGE-2 laptop, FE Sharipov performed another session with the European Neurokog experiment. Today's activities focused on virtual rotation in free floating and fixed position "corridor" passages while recording EEG (electroencephalography). [Salizhan was equipped with the Halley head electrodes. After doing the virtual turns/corridor episodes in fixed state (subject strapped down) and free-floating in zero-G called for by the Neurokog protocol, he downloaded the EEG data to a PCMCIA memory card for subsequent return to Earth, and dismantled the equipment.]

On the MedOps defibrillator equipment, successfully checked out yesterday, CDR/SO Chiao initiated the regular recharge process on battery #1 to full capacity and subsequently performed it also on battery #2. The task concluded with a battery voltage check. [Each battery was charged for ~3.5 hrs, and its open-circuit voltage was tested at the end with the Aeolus volt/amp scopemeter, then removed and stowed again. Nominally, the defib has a battery installed at all times, but with this particular unit the PDIM (power data interface module) is not functioning properly and would overcharge the batteries if left inside. They have to be charged every 60 days along with the defib checkout.]

Shuttle
Processing Status
News
Daily Mission
Return to Flight
ISS
Weekly Status
Weekly Science
Daily On-Orbit Status
Daily Crew Timeline
Soyuz | Progress
ISS News | ATV

Chiao conducted the weekly data take with the two new 17P-delivered CSA-CP units (Compound Specific Analyzer-Combustion Products), with the purpose to monitor the desired ongoing decontamination (outgassing) of the deployed instruments. Some data points were also obtained last weekend. [CSA-CP measures O2 (oxygen), CO (carbon monoxide), HCN (hydrogen cyanide), and HCl (hydrogen chloride).]

The Elektron O2 generator remains off. The FE today performed part 1 of a functional checkout on an older spare liquid unit (BZh #6), supported by tagup with ground specialists. Both micropumps were diagnosed "with bubbles". More troubleshooting is planned for Friday (4/8).

TsUP/Moscow performed an atmosphere refresh from Progress 17 oxygen storage that added 10 mmHg/Torr of O2. [The resulting ppO2 was 169.3 mmHg, with a final concentration of O2 at 22.2%, including the 6 mmHg error band on the MCA (major constituents analyzer). Flight Rule lower limit is 146 mmHg.]

With the Vozdukh carbon dioxide (CO2) removal system turned off, Salizhan replaced the second LiOH (lithium hydroxide) absorbent filter cartridges (PP) and activated it. [This is the second of 14 lifetime-expired but unused Russian LiOH canisters onboard. The two canisters are being tested for approximately 24 hours each sequentially, with the Vozdukh off for the duration of the test so that TsUP can monitor the CO2 levels to assess the efficiency of the expired LiOH. The data will be used to extrapolate efficiency of the remaining 12 LiOH canisters. After that, a determination will be made on further potential use of the units. During operation, the CO2 absorbed by the cartridges is periodically vented overboard through the BVK valve by heating (and thus regenerating) the filters.]

Leroy completed the periodic filter cleaning on the autonomous PCG-STES010 (Protein Crystal Growth-Single Locker Thermal Enclosure System) payload in the Lab. [For the PCG-STES, today is the 861st day of continuous powered operation onboard the ISS, which is unprecedented for this experiment hardware. Its previous run time onboard Mir did not exceed 200 days. Nominally, PCG-STES010 powered operations are expected to continue while plans are finalized for its return on LF-1 (STS-114) next month.]

The CDR/SO spent two hours doing the periodic thorough inspection of the ELPS (emergency lighting power sources) in the U.S. segment. [There are three ELPS units in the Node, two in the Lab, and one ELPS in the Airlock.]

The FE also worked on another Russian IPK-1M gas mask (of seven total), upgrading it with a new nozzle attachment. The work was supported "as needed" by tagup with TsUP. [This upgrading is a longer-term activity that started with the first mask on 3/10.]

Working off the voluntary Russian task list, Salizhan performed the regular daily inspection of the BIO-5 Rasteniya-2 ("Plants-2") experiment, including filling its water canister as required. [Rasteniya researches growth and development of plants under spaceflight conditions in the Lada-5 greenhouse.]

Leroy Chiao completed the routine maintenance of the SM's SOZh environment control & life support system, including ASU toilet facilities. For Salizhan, today's preparation of the regular IMS (inventory management system) "delta" file for the daily automated export/import to the three IMS databases on the ground was on his voluntary task list.

The FE also had about an hour to continue making preparations of equipment and personal items for return on Soyuz TM-5/9S.

The crew conducted their regular daily 2.5-hr. physical exercise program on TVIS treadmill, RED exerciser, and VELO cycle with bungee cord load trainer. [Salizhan's daily protocol currently prescribes a four-day microcycle exercise with 1.5 hr on the treadmill and one hour on VELO (today: Day 1 of a new set).]

Chiao then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.

At ~3:30pm EDT, the crew is scheduled for a privatized teleconference in S/G2 (Space-to-Ground 2) with the LF-1/STS-114 Shuttle crew.

For his next "Saturday Science" program on 4/9, Dr. Chiao has selected close-up photography of the SEM (Space Experiment Module) satchel and vials onboard the ISS.

Planning is underway to give the crew half a day off on 4/13 and 4/15.

The Kodak DCS 760 digital camera that failed last week and was restored by the crew over the weekend has failed again. Troubleshooting plans are in work. [This camera is designated as a Return-to-Flight item for providing photography of the STS-114/Discovery approach.]

The voltage/current scopemeter chosen yesterday for supporting the upcoming VOA (Volatile Organics Analyzer) troubleshooting was found to work only with a set of disposable C-cell batteries (instead of the nominal rechargeable units). Since the number of disposable batteries onboard is limited and the scopemeters are required for certain activities, assessment is underway to determine a potential repair for the problem of the scopemeter not charging.

Last Monday (4/4) the Russian air conditioner SKV-1 stopped functioning properly and was deactivated, either due to clogged lines or a dried wick. A troubleshooting plan is in work. Meanwhile, the SKV-2 is operating nominally to remove moisture from the air.

During the day, TsUP/Moscow conducted hardware tests on the automated passive Kurs-P rendezvous & approach radar system on the DC1 side, on both subsets (strings). [The DC1 port will be used by Soyuz 10S for docking on 4/16 (10:17pm EDT). 9S is currently docked at the FGB nadir port.]

TsUP also prepared the Progress KDU propellant system remotely for the upcoming prop transfer from both manifolds (fuel tanks BG1, BG2; oxidizer tanks BO1, BO2) to the Service Module (SM) storage tanks.

At 11:03am today, CMG-3 (control moment gyro #3) experienced a disturbance just before LOS (loss of signal). Vibration & current peaked at 0.063g and 0.7 amps, respectively, on the SMCC (Spin Motor Control Current). Upon AOS (acquisition of signal) the event had passed, and vibration & current had returned to nominal levels. Experts are looking at data retained during the LOS to determine the magnitude of the disturbance. The planned 31-hour maneuver to remove the roll/yaw bias will delayed until the signature is understood.

During a 5-min "window" today at 7:33am researchers on the ground again conducted the European/Russian laser-beam experiment SPQR (Specular Point-like Quick Reference). [SPQR, installed at SM window #3 along with its Nikon D1X camera, tests a ground-based imaging system, using special optics and image processing, to determine the feasibility of detecting external damage to a spacecraft in orbit from the ground. It uses a pyramidal corner reflector (CCR, Cube Corner Reflector) at the SM window, to reflect a laser beam emitted by a ground station back to the ground. The crew was advised not to look out the portholes during the sessions, the times of which were uplinked, and there are no CEO targets scheduled during the brief sessions. The SPQR Hazard Report indicates that the laser power at the ISS remains well below the threshold for injury.]

Today's CEO (crew earth observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, were Lahore, Pakistan (this pass approached this mega city from the NW. Although lighting and weather were expected to be excellent, surface visibilities were probably reduced by aerosols, and the crew may not have spotted the city until right over it. Looking just right of track and using the long lens to try and map the urban area), Nile River Delta (best views of this target were over the northwestern edges of the delta to the S and W of Alexandria. Careful mapping of the edges of the developed areas was desired, where changes from previous imagery should be most noticeable), and NE US Flooding (DYNAMIC EVENT: After weeks of strong storms the northeastern US is reeling from heavy rainfall and snowmelt that are causing widespread stream and river flooding. Particularly hard hit has been the state of New Jersey. As ISS tracked southeastward across this area in fair weather, the crew was to look for evidence of standing water and sediment plumes from the upper Ohio River Valley to the New Jersey Coast).

CEO photography can be viewed and studied at the websites:

See also the website "Space Station Challenge" at:

To view the latest photos taken by the expedition 10 crew visit:

Expedition 10 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Upcoming Key Events:

  • Soyuz TMA-6 (10S) launch -- 4/14 (8:46pm EDT); with Expedition 11 (CDR Sergei Krikalev, FE/SO John Phillips & VC8 cosmonaut Roberto Vittori/ESA-Italy); launch time at Baikonur: 6:46am on 4/15.
  • Soyuz TMA-6 docking -- 4/16 (10:17pm EDT);
  • Soyuz TMA-5 (9S) undocking -- 4/24 (2:38pm EDT) with Exp. 10 crew (after 193 days on orbit, 191 days on board ISS) and VC8 cosmonaut Vittori;
  • Soyuz TMA-5 landing -- 4/24 (6:04pm EDT (Kustanai: 4:04am on 4/25) ;
  • LF1 (STS-114) launch -- 5/15;
  • Progress M-53 (18P) launch -- 6/10;
  • ULF1.1 (STS-121) launch -- NET 7/12;
  • Progress M-54 (19P) launch -- 8/24;
  • Soyuz TMA-7 (11S) launch -- 9/27.

ISS Location NOW

Full Size/Update
Real Time ISS Tracker - More Links

ISS Orbit (as of this morning, 5:13am EDT [= epoch]): Mean altitude -- 356.7 km

  • Apogee height -- 361.0 km
  • Perigee height -- 352.4 km
  • Period -- 91.67 min.
  • Inclination (to Equator) -- 51.64 deg
  • Eccentricity -- 0.0006394
  • Solar Beta Angle -- -42.7 deg (magnitude decreasing)
  • Orbits per 24-hr. day -- 15.70
  • Mean altitude loss in last 24 hours -- 205m
  • Revolutions since FGB/Zarya launch (Nov. 98) -- 36440

ISS Altitude History

Apogee height -- Mean Altitude -- Perigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA's Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA's Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.